# Conversational SimulMT: Efficient Simultaneous Translation with Large Language Models

### v3

@@ Abstract
Simultaneous machine translation (SimulMT) presents a challenging trade-off between translation quality and latency. Recent studies have shown that LLMs can achieve good performance in SimulMT tasks. However, this often comes at the expense of high inference cost and latency. In this paper, we propose a conversational SimulMT framework to enhance the inference efficiency of LLM-based SimulMT through multi-turn-dialogue-based decoding where source and target chunks interleave in translation history, enabling the reuse of Key-Value cache. To adapt LLMs to the proposed conversational decoding, we create supervised fine-tuning training data by segmenting parallel sentences using an alignment tool and a novel augmentation technique to enhance generalization. Our experiments with Llama2-7b-chat on two SimulMT benchmarks demonstrate that the proposed method empowers the superiority of LLM in translation quality, meanwhile achieves comparable computational latency with specialized SimulMT models. [^1] Code, weights, and data will be released with publication.
@@ 1 Introduction
Simultaneous machine translation (SimulMT) systems provide real-time translation of text input stream Gu et al. (2017). This task plays an important role in real-world applications, such as facilitating communication in online conferences and generating live subtitles with strict latency requirements. In online speech translation, where re-translation is impossible, SimulMT should be robust to errors propagated from the automatic speech recognition system, and produce high-quality translation with low latency.
(a) Offline Prompt
(b) Conversational Prompt
Figure 1: SimulMT prompts in Llama2 format. The system message is blue. The source and corresponding target translations chunks have matching color codes.
@@
Although large language models (LLMs) have shown the potentials in machine translation (Hendy et al., 2023; Zhu et al., 2023), their applications to SimulMT is non-trivial, as they are not inherently designed for simultaneous decoding. Recent works have attempted to adapt LLMs for SimulMT with prefix fine-tuning, incremental decoding (Wang et al., 2023b) and learning to wait for more source tokens before translation (Koshkin et al., 2024). These works show LLMs, with careful prompt-engineering, could approach the performance of specialized SimulMT models. However, high computational cost, slow inference, and high latency render these approaches impractical for real-world applications (Yuan et al., 2024). This is primarily due to the use of offline prompting, where arriving source tokens are inserted at the end of the source sequence, disrupting the continuity of the translation history (Figure 1(a)). This prevents reusing cached target history states and requires re-computation of source and target representations.
@@
To mitigate this issue, we propose conversational prompts that resemble the multi-turn dialogue nature of LLMs. Specifically, user inputs are treated as the source tokens to be read, while the LLM’s responses are considered the predicted target tokens to be written. In our conversational SimulMT, newly arrived source form the current instruction, while previous source tokens and their translation are treated as conversation history (Figure 1(b)). This conversational prompt format enables the reuse of Key-Value cache Pope et al. (2023), as all content is appended incrementally without modifying the translation history, hence improving inference efficiency. However, conversational SimulMT poses a new challenge for LLMs to accurately identify source sequences, particularly when these sequences are interrupted with target chunks within the conversation history.
@@
Interleaving incomplete source and target segments in the dialogue history is unnatural (see Figure 1(b)). The code-switching phenomenon is exhibited to some extent in LLMs for some languages (Yong et al., 2023); however, it is the continuation rather than the translation of the previous content, making it challenging to leverage existing code-switched datasets for training. To adapt the LLM to the conversational decoding format, we take the approach of supervised fine-tuning (SFT) in the conversational SimulMT. The challenge is the lack of the conversational SimulMT data for SFT. We generate the training data by segmenting parallel sentence pairs into smaller chunks based on a transformation of the word alignments. The segmented chunks are further augmented to handle different latency requirements in SimulMT.
@@
Experiments on two SimulMT benchmarks demonstrate the effectiveness of our proposed conversational SimulMT in balancing the trade-offs between accuracy, speed and flexibility to different latency requirements. Compared to offline prompting, our method not only maintains strong performance, but also benefits from reduced latency. Notably, our method attains similar decoding speed to the offline LLM-based MT.
@@
  • We introduce the conversation prompt to reduce the inference cost of LLM-based SimulMT by leveraging its multi-turn dialogue capability and enabling efficient reuse of Key-Value cached computations.
  • We present an automated training data curation pipeline which involves a simple heuristic and a novel augmentation strategy to efficiently extract source and target chunks from their word alignments, enabling the LLM to decode with prefix-growing code-switched inputs to support processing of incoming source words while generating target words.
  • Experiments demonstrate that the proposed conversational SimulMT obtains up to 2$\times$ acceleration compared to the offline-prompting baseline while maintaining comparable translation quality, emphasizing its value in practical applications.
@@ 2 Background
Unlike offline machine translation, where models generate target translation $\mathbf{y}=(y_{1},...,y_{J})$ given a complete source sentence $\mathbf{x}=(x_{1},...,x_{I})$, SimulMT incrementally translates in an online manner with partial source context $\mathbf{x}_{\leq t}=(x_{1},...,x_{t})$ where $I$ and $J$ are source and target length, and $t\leq I$. A core component of SimulMT is a read-write policy to decide at every step whether to wait for one or multiple new source tokens (READ) or to generate target tokens given the current partial source sentence (WRITE). This policy is crucial in balancing the trade-offs between translation quality and latency.
@@
Studies have explored equipping the capability of simultaneous decoding to an offline neural machine translation model by repetitively performing offline decoding on the incrementally updated source and target histories, without the necessity of READ -WRITE policy learning and model customization (Liu et al., 2020; Nguyen et al., 2021; Polák et al., 2022; Guo et al., 2023). Specifically, it involves a chunk-wise READ policy that reads a fixed number of $n$ tokens in each round and a WRITE policy that selects the longest common prefix (LCP) among the beam candidates as a stable hypothesis to be committed Polák et al. (2022). LCP often results in high latency when no candidates share the same initial token, hindering the WRITE action. Relaxed Agreement Longest Common Prefix (RALCP) (Wang et al., 2023b) employs a voting mechanism to accept a common prefix if the agreement among candidates is above an agreement threshold $\gamma$. Once a hypothesis is committed, it is no longer modified.
@@
Given that incremental decoding essentially repeats performing offline decoding, leveraging offline-style translation prompts for LLMs is a straightforward approach. Meanwhile, offline-style prompts align with LLMs’ innate ability to follow instructions (translating content from one language to another), avoiding the additional fine-tuning Xu et al. (2023). During each incremental decoding round, a new source chunk is READ in and appended to the source history $\mathbf{x}$. The translation hypothesis is WRITTEN out and appended to target history $\mathbf{y}$, forming the latest incremental states. Given the new source and target history $\mathbf{x}$ and $\mathbf{y}$, LLMs are prompted to generate the latest translation with the offline prompt as Figure 1(a).
@@ 3 Conversational SimulMT
While incremental decoding with offline prompt enables LLMs to perform simultaneous decoding, it faces high computational latency due to the insertion of newly arrived source tokens in the middle of prompt, disrupting the reuse of cached target history states. In this section, we propose conversational prompts to improve the decoding efficiency and balance quality-latency trade-off.
@@ 3.1 Decoding with Conversational Prompt
The efficiency improvement in LLMs hinges on maintaining the Key-Value (KV-) cache reuse, i.e. the decoding process must consistently add new tokens at the end of the sequence without altering the middle elements. When LLMs are performing multi-turn dialogues, the prompt for each turn is composed of a user input and agent response, separated by [INST] and [/INST] tags: “<s>[INST] $X$ [/INST] $Y$</s>", and conversation histories are simply concatenated as the context Touvron et al. (2023). Drawing parallels to multi-turn dialogues in LLMs, SimulMT can also be viewed similarly, where user inputs are equivalent of READ action, and agent responses are equivalent of WRITE action. At round $t$, LLM reads a source context chunk $X_{t}$ and writes its translation $Y_{t}$: “<s>[INST] $X_{t}$ [/INST] $Y_{t}$</s>". The conversational prompt of SimulMT can be constructed by placing the source chunk as user input, and the model produces the translation as a response (see Figure 1(b)). The conversational prompt enables re-using KV-cache as all contents are appended incrementally. Our approach is also able to leverage the hypothesis selection strategy of RALCP (Wang et al., 2023b) to prune the unstable suffixes in each response. Algorithm 1 in Appendix B presents the detailed decoding process.
@@
Table 1: Performance of zero-shot and few-shot conversational SimulMT (Convers.) with Llama2-7b-chat on the WMT15 De->En test set. Offline decoding result with the same model is presented (Offline MT) as the baseline. A failure case is also shown, in which the LLM tends to complete the sentence in the translation with hallucination (red) leading to a more serious failure in the second chunk.
We conduct a pilot experiment on the zero-shot and few-shot capability of LLMs in decoding with conversation prompts. Specifically, we evaluate the pretrained Llama2-7b-chat (Touvron et al., 2023) with zero- and five-shot settings on the WMT15 De->En test set with chunk size $n$ set as 5. As illustrated by Table 1, SimulMT with conversational prompts obtain inferior performance even when instructed with 5-shot prompt. The failure case example indicates that the LLM cannot accurately translate the content within the partial source sentence. Instead, it tends to complete the sentence in the translation with hallucinated content. This is because the LLM is trained on full sentences and, therefore, is not able to handle interleaved partial sentence translation in the multi-turn dialogue format. To this end, we propose to perform supervised fine-tuning (SFT) with the LLM on SimulMT data in a conversational format. Our approach to tackle the challenge of lacking the conversational SimulMT data for SFT is introduced in the following section.
Figure 2: The illustration of the data curating process. The first graph is obtained from fast_align, it is then modified into a monotonic dependency graph by adding additional edges. The Meta Trajectory can be derived by segmenting the monotonic dependency graph with minimal dependency (segment with the dotted line in graph 3). Finally, Policy Generalization is applied to augment the segmented graph with merge (red dotted lines will be removed) and shift (blue dotted lines are shifted) operations. Chunks in the trajectories derived from the third and fourth graphs are highlighted with different colors.
@@ 3.2 SFT on Conversational SimulMT Data
As conversational SimulMT data is not naturally available, we propose to synthesize READ / WRITE chunks by segmenting sentence pairs from parallel corpora. Inspired by Arthur et al. (2021) which generates the oracle policy from word alignments, we further extend the approach by carefully addressing the impact of word reordering and improving the generalizability of the oracle policy. Specifically, we first build monotonic dependency graph from the alignment of a sentence pair. We then segment the graph and convert these segments into READ / WRITE pairs, followed by augmentation to improve its generalization across various latency demands (Figure 2). The process is explained below.
@@
Given a sentence pair, we employ fastalign Dyer et al. (2013) to obtain word alignment between source and target tokens (Step 1 in Figure 2). The obtained alignment is a set $\mathcal{A}$ of pairs $(i,j)$ denoting the source token $x_{i}$ is aligned with its corresponding target token $y_{j}$. We define the sufficient source token set to generate a given target token $y_{j}$ as $\textbf{a}_{j}=\{i|(i,j)\in\mathcal{A},\forall i\in[0,I]\}$ of all source tokens needed to produce the target token $y_{j}$.
A source and target sentences have a monotonic translation relationship if the previous target tokens only aligned with the previous source tokens, i.e. $\forall j>k\quad\max(\textbf{a}_{j})\geq\max(\textbf{a}_{k})$ (Koehn et al., 2005; Ling et al., 2011). This condition ensures that the relative order of words is preserved between the source and target sentences. In that case, the optimal minimum-latency policy that retains sufficient source information is to produce the monotonic translation that follows the word order of the source sequence, i.e. WRITE the target token $y_{j}$ after reading the final source token $x_{\max(\textbf{a}_{j})}$ needed for its translation, and READ the next source tokens.
@@
The monotonic dependency between source and target sentences enables the effective implementation of an optimal READ /WRITE policy. However, in practice, translations often exhibit non-monotonic relationships, which require re-ordering to produce grammatically correct output in the target language, especially when the source and target languages have significantly different syntactic structures. To address this, we propose a heuristic method for constructing a monotonic dependency graph $\overrightarrow{\mathcal{A}}$ from the alignment set $\mathcal{A}$, such that the monotonic condition is met (Step 2 in Figure 2).
More specifically, for each target token $y_{j}$ which does not satisfy the monotonic condition $\max(\textbf{a}_{j})<\max(\textbf{a}_{j-1})$, we add a new edge from the last sufficient source token $x_{\max(\textbf{a}_{j-1})}$ to $y_{j}$ to eliminate the need of re-ordering. For example, in Figure 2, $y_{2}$ violates the monotonic condition as the last source token in its sufficient set $\textbf{a}_{2}=\{1\}$ comes after the last sufficient source token of the previous target $\textbf{a}_{1}=\{1,2\}$. Therefore, a new edge from $x_{2}$ to $y_{2}$ is added to the graph.
@@
We then segment the monotonic dependency graph and convert these segments into READ / WRITE pairs, representing the meta trajectory of the oracle policy with minimum latency (Step 3 in Figure 2). This segmentation involves examining each target token and identifying its exclusive corresponding source tokens with minimal dependency. Each subgraph $\overrightarrow{\mathcal{A}_{j}}$ corresponds with a READ / WRITE pair ($R_{j}$, $W_{j}$) where $W_{j}=\{y_{j}\}$ can be a single target token and $R_{j}=\{x_{i}|i\in\textbf{a}_{j}\setminus\textbf{a}_{j-1}\}$ is a set of new source tokens since the last target token. When multiple consecutive target tokens depend on the same source token, we combine their WRITE actions and assign the shared source token to $R_{j}=\{x_{i}\}$, forming $W_{j}=\{y_{j},...,y_{j+n}\}$ with these target tokens. This procedure generates a meta READ / WRITE trajectory for the oracle policy, denoted as $RW^{\star}=[(R_{1},W_{1}),...,(R_{C},W_{C})],C\leq I$, representing $C$ chunks.
@@
Since the meta trajectories are tailored for minimal latency, they may not generalize well to different lengths of the input chunk, corresponding to different levels of latency. To improve the LLM’s adaptability across various latency demands, we augment the meta-trajectory $RW^{\star}$ with a series of merge and shift operations (Step 4 in Figure 2). We first traverse $RW^{\star}$ and randomly merge $\delta$ consecutive READ and WRITE actions, forming new pairs $([R_{c},...,R_{c+\delta}],[W_{c},...,W_{c+\delta}])$, where []delimited-[][$\cdot$][ ⋅ ] is the string concatenation operation. Here, $\delta$ is a variable re-sampled from a uniform distribution $\mathcal{U}(\delta_{\min},\delta_{\max})$ where $\delta_{\min}$ and $\delta_{\max}$ are predefined hyperparameters.
Additionally, with a probability of $\beta$, we randomly shift a portion of tokens from a WRITE action $W_{c}$ to the next one $W_{c+1}$ in the merged trajectory. More specifically, we split $W_{c}$ at a proportion $\rho$ and transfer the latter part to the next pair, resulting in $(R_{c},W_{c}^{<\rho}),(R_{c+1},[W_{c}^{>\rho},W_{c+1}])$ where $\rho$ is sampled from $\mathcal{U}(\rho_{\min},0.9)$ where $\rho_{\min}$ is a hyperparameter. The proposed augmentation strategy not only enhances the LLM’s context conditioning ability but also suits the incremental decoding, where the end of the prediction is often truncated by the stable hypothesis selection algorithm.
The resulting trajectory is a sequence of READ / WRITE chunks where each READ chunk contains source text of differing lengths. We then format them in the conversational format to make the trajectory into a text sequence. During SFT, we apply cross-entropy loss only on target tokens within WRITE chunks without the shift operation.

### v4

@@ Abstract
Simultaneous machine translation (SimulMT) presents a challenging trade-off between translation quality and latency. Recent studies have shown that LLMs can achieve good performance in SimulMT tasks. However, this often comes at the expense of high inference costs and latency. In this paper, we propose a conversational SimulMT framework to enhance the inference efficiency of LLM-based SimulMT through multi-turn-dialogue-based decoding where source and target chunks interleave in translation history, enabling the reuse of Key-Value cache. To adapt LLMs to the proposed conversational decoding, we create supervised fine-tuning training data by segmenting parallel sentences using an alignment tool and a novel augmentation technique to enhance generalization. Our experiments with Llama2-7b-chat on three SimulMT benchmarks demonstrate that the proposed method empowers the superiority of LLM in translation quality, meanwhile achieving comparable computational latency with specialized SimulMT models. [^1] Our code is available at https://github.com/yuriak/LLM-SimulMT
@@ 1 Introduction
Simultaneous machine translation (SimulMT) systems provide real-time translation of text input stream gu-etal-2017-learning. This task plays an important role in real-world applications, such as facilitating communication in online conferences and generating live subtitles with strict latency requirements.
Figure 1: Comparison of offline prompt (left) and conversational prompt (right). Offline prompt inserts tokens mid-sequence, preventing KV-cache reuse (red X), while conversational prompt appends content sequentially, enabling efficient cache utilization (blue blocks).
@@
Although large language models (LLMs) have shown the potentials in machine translation (DBLP:journals/corr/abs-2302-09210; DBLP:journals/corr/abs-2304-04675), their applications to SimulMT is non-trivial, as they are not inherently designed for simultaneous decoding. Recent works have attempted to adapt LLMs for SimulMT with prefix fine-tuning, incremental decoding (wang2023simultaneous) and learning to wait for more source tokens before translation (koshkin2024transllama). These works show LLMs, with careful prompt-engineering, could approach the performance of specialized SimulMT models. However, high computational cost, slow inference, and high latency render these approaches impractical for real-world applications (yuan2024llm). This is primarily due to the use of offline prompting, where arriving source tokens are inserted at the end of the source sequence, disrupting the continuity of the translation history (Figure˜1 left). This prevents reusing cached target history states and requires re-computation of source and target representations.
@@
To mitigate this issue, we propose conversational prompt that resemble the multi-turn dialogue nature of LLMs. Specifically, user inputs are treated as the source tokens to be read, while the LLM’s responses are considered the predicted target tokens to be written. In our conversational SimulMT, newly arrived source form the current instruction, while previous source tokens and their translations are treated as conversation history (Figure˜1 right). This conversational prompt enables the reuse of Key-Value cache kv-cache2023, as all content is appended incrementally without modifying the translation history. However, conversational SimulMT poses new challenges for LLMs to comprehend the segmented source content and produce a coherent translation via multi-turn conversation.
@@
To adapt the LLM to the conversational decoding format, we opt to perform supervised fine-tuning (SFT) on the pretrained LLM. But the challenge is the lack of the conversational SimulMT data for SFT. Interleaving incomplete source and target segments in the dialogue history is unnatural (see Figure˜1). This code-switching style is exhibited in some languages (yong2023prompting); however, it is the continuation rather than the translation of the previous content, making it challenging to leverage existing code-switched datasets for training. Therefore, we propose to curate the training data by segmenting parallel sentence pairs into smaller chunks based on a transformation of the word alignments. The segmented chunks are further augmented to handle different latency requirements.
@@
Experiments on three SimulMT benchmarks demonstrate the effectiveness of our proposed conversational SimulMT in balancing the trade-offs between accuracy, speed and flexibility to different latency requirements. Compared to offline prompting, our method not only maintains strong performance, but also benefits from reduced latency. Notably, our method attains similar decoding speed to the LLM-based OfflineMT.
@@
  • We introduce conversational prompting to reduce the inference cost of LLM-based SimulMT by leveraging its multi-turn dialogue capability and enabling efficient reuse of Key-Value cached computations.
  • We present an automated training data curation pipeline that can turn any offline translation parallel corpus into the conversational prompt format and generalize with a novel augmentation strategy into any inference setting.
[…]
@@ 2 Background
Unlike offline machine translation (OfflineMT), where models generate target translation $\mathbf{y}=(y_{1},...,y_{J})$ given a complete source sentence $\mathbf{x}=(x_{1},...,x_{I})$, SimulMT incrementally translates with partial source context $\mathbf{x}_{\leq t}=(x_{1},...,x_{t})$ where $t\leq I$. A core component of SimulMT is a read-write policy that decides whether to wait for new source tokens (READ) or generate target tokens (WRITE), balancing translation quality and latency.
@@
Studies have explored adapting OfflineMT models for simultaneous decoding by performing offline decoding on incrementally updated histories (low_latency; nguyen21c_interspeech; polak-etal-2022-cuni; HW_TSC). This involves a chunk-wise READ policy that reads $n$ tokens per round and a WRITE policy that commits stable partial translations using the longest common prefix (LCP) polak-etal-2022-cuni algorithm. LCP often causes high latency when candidates lack common initial tokens. Relaxed Agreement LCP (RALCP) (wang2023simultaneous) was proposed to vote for accepting prefixes with candidate agreement above threshold $\gamma$.
@@
Since incremental decoding essentially repeats offline decoding, using offline-style translation prompts with LLMs is straightforward and aligns with their instruction-following capabilities xu2023paradigmshift. During each round, a source chunk is READ and appended to source history $\mathbf{x}$. LLMs generate translations using offline prompts as shown in Figure˜1, which are then WRITTEN to target history $\mathbf{y}$.
@@ 3 Conversational SimulMT
While incremental decoding with offline prompt enables LLMs to perform simultaneous decoding, it faces high computational latency due to the insertion of newly arrived source tokens in the middle of the prompt, disrupting the reuse of cached target history states. In this section, we propose conversational prompts to improve the decoding efficiency and balance quality-latency trade-off.
@@ 3.1 Decoding with Conversational Prompt
The efficiency improvement in LLMs hinges on maintaining the Key-Value (KV-) cache reuse, i.e. the decoding process must consistently add new tokens at the end of the sequence without altering the middle elements. When LLMs are performing multi-turn dialogues, the prompt for each turn is composed of a user input and assistant response separated by special tokens, and conversation histories are simply concatenated as the context touvron2023llama2. Drawing parallels to multi-turn dialogues in LLMs, SimulMT can also be viewed similarly, where user inputs and assistant responses are equivalent to READ and WRITE action. At round $t$, LLM reads a source context chunk $X_{t}$ and writes its translation $Y_{t}$: “<U> $X_{t}$ <A> $Y_{t}$". The already processed chunks are concatenated as contexts, serving the latest translation round of new incoming chunks. As all contents are appended incrementally, the reuse of KV-cache becomes feasible again like in multi-turn dialogue (see Figure˜1). Our approach also adapts the hypothesis selection strategy e.g. RALCP (wang2023simultaneous) to prune the unstable suffixes in each response. Algorithm˜1 in Appendix˜A presents the detailed decoding process.
@@
Table 1: Performance comparison of Llama2-7b-chat on WMT15 De->En test set in zero-shot and few-shot conversational SimulMT settings. OfflineMT results are included as a baseline. The example failure case demonstrates how the LLM hallucinates completions (shown in red) when translating partial sentences, leading to compounding errors in subsequent chunks.
We conducted a pilot experiment to assess LLMs’ zero-shot and few-shot capabilities with conversational prompts. Using Llama2-7b-chat (touvron2023llama2) on the WMT15 De->En test set with chunk size $n=5$, we tested both zero- and five-shot settings. As shown in Table 1, conversational SimulMT performed poorly even with 5-shot prompting. The failure analysis reveals that LLMs, trained primarily on complete sentences, struggle with partial source translation and tend to hallucinate completions when presented with fragments in a multi-turn dialogue format. To address this limitation, we propose to SFT LLMs on conversational SimulMT data. The following section details our approach to converting a normal bi-text corpus into conversational prompt format.
Figure 2: The illustration of the data curating process. The first graph is obtained from fast_align, it is then modified into a monotonic dependency graph by adding additional edges. The Meta Trajectory can be derived by segmenting the monotonic dependency graph with minimal dependency (segment with the colored solid line in step 3). Finally, Policy Generalization is applied to augment the segmented graph with merge (red dotted lines will be removed) and shift (blue dotted lines are shifted) operations. Chunks in the trajectories derived from the third and fourth graphs are highlighted with different colors.
@@ 3.2 SFT on Conversational SimulMT Data
As conversational SimulMT data is not naturally available, we propose to synthesize READ / WRITE chunks by segmenting sentence pairs from parallel corpora. Inspired by arthur-etal-2021-learning which generates the oracle policy from word alignments, we further extend the approach by carefully addressing the impact of word reordering and improving the generalizability of the oracle policy. Specifically, we first build monotonic dependency graph from the alignment of a sentence pair. We then segment the graph and convert these segments into READ / WRITE pairs, followed by augmentation to improve its generalization across various latency demands (Figure˜2). The process is explained below.
@@
Given a sentence pair, we employ fastalign fastalign to obtain word alignment between source and target tokens (Step 1 in Figure˜2). The obtained alignment is a set $\mathcal{A}$ of pairs $(i,j)$ denoting the source token $x_{i}$ is aligned with its corresponding target token $y_{j}$. We define the sufficient source token set to generate a given target token $y_{j}$ as $\textbf{a}_{j}=\{i|(i,j)\in\mathcal{A},\forall i\in[0,I]\}$.
A source and target sentences have a monotonic translation relationship if the previous target tokens only aligned with the previous source tokens, i.e. $\forall j>k\quad\min(\textbf{a}_{j})\geq\max(\textbf{a}_{k})$ (koehn-etal-2005-edinburgh; ling-etal-2011-discriminative). This condition ensures that the relative order of words is preserved between the source and target sentences. In that case, the optimal minimum-latency policy that retains sufficient source information is to produce the monotonic translation that follows the word order of the source sequence, i.e. WRITE target token $y_{j}$ immediately after reading the final required source token $x_{\max(\textbf{a}_{j})}$, and then READ the next source tokens.
@@
Monotonic dependency enables effective implementation of optimal READ /WRITE policies. However, translations often require reordering to produce grammatically correct output, especially between languages with different syntactic structures. To address this, we propose constructing a monotonic dependency graph $\overrightarrow{\mathcal{A}}$ from alignment set $\mathcal{A}$ (Step 2 in Figure˜2) such that the monotonic condition is met.
For each target token $y_{j}$ violating the monotonic condition $\min(\textbf{a}_{j})<\max(\textbf{a}_{j-1})$, we add a new edge from the last sufficient source token $x_{\max(\textbf{a}_{j-1})}$ to $y_{j}$, eliminating the need for reordering. In Figure˜2, $y_{2}$ violates monotonicity as its earliest required source token $\min(\textbf{a}_{2})=1$ precedes the last required source token for the previous target $\max(\textbf{a}_{1})=2$. Thus, we add an edge from $x_{2}$ to $y_{2}$.
@@
We then segment the monotonic dependency graph and convert these segments into READ / WRITE pairs, representing the meta trajectory of the oracle policy with minimum latency (Step 3 in Figure 2). We examine each target token to identify its exclusive corresponding source tokens with minimal dependency. Each subgraph $\overrightarrow{\mathcal{A}_{j}}$ corresponds to a pair $(R_{j},W_{j})$ where $W_{j}=\{y_{j}\}$ is a target token and $R_{j}=\{x_{i}|i\in\textbf{a}_{j}\setminus\textbf{a}_{j-1}\}$ contains new source tokens required since the previous target. When consecutive target tokens depend on the same source token, we combine their WRITE actions, assigning the shared source token to $R_{j}=\{x_{i}\}$ and forming $W_{j}=\{y_{j},...,y_{j+n}\}$. This generates a meta trajectory $RW^{\star}=[(R_{1},W_{1}),...,(R_{C},W_{C})],C\leq I$, with $C$ chunks.
@@
Since the meta trajectories are tailored for minimal latency, they may not generalize well to different lengths of the input chunk, corresponding to different levels of latency. To improve the LLM’s adaptability across various latency demands, we augment the meta-trajectory $RW^{\star}$ with a series of merge and shift operations (Step 4 in Figure 2). We first traverse $RW^{\star}$ and randomly merge $\delta$ consecutive READ and WRITE actions, forming new pairs $([R_{c},...,R_{c+\delta}],[W_{c},...,W_{c+\delta}])$, where $\cdot$] is the string concatenation operation. Here, $\delta$ is a variable re-sampled from a uniform distribution $\mathcal{U}(\delta_{\min},\delta_{\max})$ where $\delta_{\min}$ and $\delta_{\max}$ are predefined hyperparameters.
Additionally, with a probability of $\beta$, we shift a portion of tokens from a WRITE action $W_{c}$ to the next one $W_{c+1}$ in the merged trajectory. More specifically, we split $W_{c}$ at a proportion $\rho$ and transfer the latter part to the next pair, resulting in $(R_{c},W_{c}^{<\rho}),(R_{c+1},[W_{c}^{>\rho},W_{c+1}])$ where $\rho$ is sampled from $\mathcal{U}(\rho_{\min},0.9)$ where $\rho_{\min}$ is a hyperparameter.
This augmentation enhances the LLM’s context conditioning and suits incremental decoding where prediction endings are often truncated by hypothesis selection algorithms. The resulting trajectory consists of READ /WRITE chunks of varying lengths, formatted with conversational prompts for SFT. During training, we apply cross-entropy loss only on target tokens within unshifted WRITE chunks.